The badminton court shown below occupies 1,456 square feet of the floor space of a gymnasium. If its length is 4 feet more than twice its width, find its dimensions.
width= ft
length= ft
The dimensions i was given on rectangle are length is <---x--->, and the width is <---2x+4--->

Answer by stanbon(75887)

The badminton court shown below occupies 1,456 square feet of the floor space of a gymnasium. If its length is 4 feet more than twice its width, find its dimensions.
width= ft
length= ft
The dimensions i was given on rectangle are length is <---x--->, and the width is <---2x+4--->
=============
Equation:
x(2x+4) = 1456
-----
2x^2 + 4x - 1456 = 0
-----
x^2 + 2x - 728 = 0
------
(x-26)(x+28) = 0
------------------------
Positive solution:
x = 26 ft
2x+4 = 56 ft
------------
Cheers,
Stan H>
-------------